# Identification of three new ULXs in NGC 4631 and NGC 1097: evidence for stellar-mass black holes

### v1

@@ 3.2 Short- and long-term X-ray variability
To investigate short-term X-ray variability of the three newly identified ULXs, light curves were extracted for all available Chandra data in the 0.5–8 keV band, while the 0.3–10 keV band was used for the XMM-Newton observations. The analysis was performed using time bins of 100, 500, and 1000 seconds. For this, the CIAO and SAS tools dmextract and evselect were utilized to background-extract light curves. For the absolute timing analysis, we applied the axbary task to the event file to correct the photon arrival times to the solar system barycenter, thereby accounting for delays introduced by the orbital motion of the Earth and the Chandra spacecraft.
The Lomb-Scargle periodogram (LS) and Fast Fourier Transform (FFT) methods were applied to the light curves to search for periodic signals or variability. Only for NGC 4631 ULX X-6 light curve (see panel a of Figure 4), in the C2 observation, A periodic variation at P = 5120 s ( 1.42 hr) was detected with a 2.6- significance level, it was identified using both the LS method and with XRONOS from the power density spectrum (seepanel b of Figure 4). Moreover, we present the folded light curve for this period in panel c of Figure 4). Furthermore, long-term light curves of X-6 and X-7 were constructed using available Chandra data to study LX variability. These luminosities were derived from the best-fit reduced C-stat values obtained in the spectral analysis (Tables 1 and 2). For NGC 1097 ULX-3, a long-term X-ray light curve was constructed using all available archival Swift/XRT PC mode observations, in addition to data from Chandra and XMM-Newton. The Swift/XRT light curve was generated using a custom Python script (swifttools) designed for Swift/XRT analysis. The source and background regions were defined as circular regions with radii of 30′′ and 60′′, respectively. The resulting count rates were converted to LX based on spectral model parameters, given in the Table 3, by using WebPIMMS tool within the 0.3-10 keV energy range. For XMM-Newton, luminosity was derived from MOS data based on our spectral modeling, while for Chandra observations, we utilized the srcflux tool. The light-curve of ULX-3 is shown in Figure 6 (left panel).
@@ 3.3 Spectral evolution and state transition
We investigated the spectral evolution of NGC 4631 X-6 and X-7 using Chandra observations, excluding non-detections. Soft (S), medium (M), and hard (H) bands were defined as 0.3–1.2 keV, 1.2–2 keV, and 2–8 keV, respectively, and the M/S and H/M hardness ratios (HR) were calculated. The temporal evolution of the hardness ratios is shown in left panel of Figure 5, and the hardness–intensity diagram is presented in right panel of Figure 5. In cases where any band did not have enough counts, we adopted 3 upper limits. In addition, to investigate the potential spectral state transitions of NGC 1097 ULX-3, we constructed a hardness–intensity diagram (right panel of Figure 6), using count rates in the 0.3–1.5 keV (S) and 1.5–10 keV (H) bands. Furthermore, the count rates from Chandra and XMM-Newton observations were normalized to those from Swift/XRT using WebPIMMS to account for differences in instrumental responses.
@@ 3.4 Optical and infrared photometry
For all drizzled HST/WFC3/UVIS and HST/ACS/WFC images, point-like sources were identified using the daofind task, and their fluxes were measured with aperture photometry employing the DAOPHOT package (Stetson, 1987) within IRAF222https://iraf-community.github.io/ (Image Reduction and Analysis Facility). For the JWST/NIRCam observations, background estimation, source detection, and photometry were performed following the methodology of Allak (2023). Instrumental magnitudes were converted into Vega magnitudes using the zero points (ZP) provided for WFC3/UVIS and ACS/WFC taken from Calamida et al. (2022), and the ZP calculator333https://acszeropoints.stsci.edu/. Sources were required to be detected at a significance above 3 relative to the local background. Photometry was then measured with circular apertures of 3 pixels in radius, with the background contribution subtracted using an annulus located nine pixels from the centroid.
Using the astrometric corrections reported by Guo et al. (2023), we identified three (X6a-Xc) and two (X7a and X7b) optical counterparts for X-6 and X-7, respectively, in the F606W and F814W bands. For NGC 1097 ULX-3, we searched for its optical and NIR counterparts using available catalogs (2MASS and Gaia DR3). Aside from the galaxy nucleus, no reference sources with XRBs were identified in these catalogs; therefore, the nucleus was used as the reference source for the astrometric corrections. We corrected the relative astrometry between Chandra (CH7) and HST (F555W), Chandra and JWST (F115W), and HST and JWST, obtaining confidence error radii of , , and , respectively. Based on these results, we identified a unique optical and NIR counterpart within the astrometric error radius by applying a 3 detection threshold over the background. The positions of the ULXs are shown in Figure 7 on the Pan-STARRS RGB composite images (i, r, g bands), illustrating their distribution across the galaxy. In the NIRCam images (see Figure 8), several faint NIR sources are present around the counterpart, but they were not detected above the 1.5 level. Therefore, we adopted the optical counterpart as the corresponding NIR counterpart of ULX-3. The positions of the counterparts of three ULXs are shown in Figure 8. The photometric results of the optical and NIR counterparts of NGC 4631 ULXs X-6 and X-7, and NGC 1097 ULX-3 are presented in Tables 4 and 5.
@@
The CMDs can offer insights into the ages of the stellar populations surrounding ULXs, which may provide clues about the nature of their donor stars. PARSEC (Bressan et al., 2012) isochrones models from widely used stellar evolution codes are available for various telescope filters, including those of HST/ACS. We derived upper limit extinction, AV, using the method of Guo et al. (2023) and used this value to correct the isochrones. Using the average NH values given in Tables 1 and 2, E(F606W-F814W) values of 2.90 mag and 1.72 mag were derived for X-6 and X-7, respectively. The CMDs, plotted as F606W vs. F606W-F814W, are shown for X-6 and X-7 in Figure 9, including isochrones and field sources located within 2′′ radius circles centered on the ULX positions, alongside the optical counterparts. For ULX-3, only a near-infrared CMD could be constructed due to the limitations of the optical data. The HST/F814W photometry shows variability likely originating from the accretion disk, and the non-simultaneous filters prevent reliable color measurements. The source is extremely faint in the UVIS bands, and the F438W detection is consistent with an upper limit, making an optical CMD infeasible. In contrast, the JWST/NIRCam data show no variability, allowing a single-epoch NIR CMD to be constructed. The near-infrared CMD was constructed using F115W and F277W photometry, adopting a distance of 16.8 Mpc and an upper limit extinction of 0.06 mag (see Figure 10).
For ULX-3, the observed magnitudes were converted into fluxes using the corresponding zeropoints, and the pivot wavelength of each JWST/NIRCam filter was adopted. The resulting SED is well-fitted by a blackbody model with a best-fit temperature of T = 3300 140 K, and a reduced chi–square of = 0.95 for four degrees of freedom (see Figure 11).
4 Results and discussions
@@ 4.0.1 Spectral modeling and variability
In NGC 4631, two new transient ULX candidates, X-6 and X-7, were identified from observations conducted between 2021 and 2023. Across 13 Chandra observations (C1–C13) and three XMM-Newton observations (XM1–XM3), X-6 was not detected in C1, XM1, and XM3. In the case of X-7, the source was detected only in XM2 and in five Chandra observations (C2–C7). Furthermore, to investigate the visibility of transient ULX X-6 and X-7, we also examined Swift/XRT PC observations using the Swift-XRT data products tool444https://www.swift.ac.uk/user_objects/
@@
For X-6, both the power-law and diskbb models yield values that vary by nearly an order of magnitude in Chandra observations when intrinsic absorption is left free. The variability may reflect genuine changes in the local environment, although it could also result from statistical uncertainties or limitations of the spectral modeling. Consequently, we fixed to its average value and repeated the spectral fits (see Table 1). After applying this procedure, notable improvements in the reduced values were obtained for both the diskbb and power-law models. When applying simple phenomenological models to the X-6 spectra, we noticed that the two models could not be distinguished in Chandra datasets, probably due to limited data quality. The power-law photon index () was found to range between 1.6 and 2.8. Typically, for Galactic BH binaries (BHBs), in the range 1.4–2.4 typically correspond to the hard state, characterized by Comptonization of disk photons in a hot and thin optically thin corona (e.g., Remillard and McClintock, 2006; Done et al., 2007). For X-6, the remains ¡2.4 in nearly all Chandra observations (except C8), indicating that the source resides predominantly in the hard state. The X-6 exhibited LX in the 0.5–8 keV band of erg s-1 as derived from the power-law model.
@@
Furthermore, X-6 is well described by the diskbb model, with the inner disk temperature () varying between 0.8 and 2 keV. From diskbb spectral model, X-6 exhibited LX of erg s-1 in the 0.5–8 keV band. In the case of X-7, among all the tested single and multi-component models in XSPEC, only the diskbb model provided acceptable fits, yielding inner disk temperatures of – keV for approximately half of the available dataset. For the C2–C7 observations, X-7 exhibited LX in the 0.5–8 keV band of erg s-1 as derived from the diskbb spectral model. The temperatures indicate that X-ray emission is dominated by accretion disk processes, consistent with Galactic XRBs in the thermal state at high accretion rates. Such disk-dominated spectral behavior has been reported in several ULX studies, where thermal components consistent with super-Eddington accretion or slim disk models have been identified (e.g., Gladstone et al., 2009; Luangtip et al., 2016). Moreover, to search for possible spectral state transitions, we constructed hardness ratios as a function of time and hardness–intensity diagrams (Figure 5). Source X-6 predominantly occupied the hard state, with episodes indicating transitions from hard to intermediate and back to hard, whereas source X-7 was mostly found in the intermediate state.
@@
ULXs X-6 and X-7 do not show significant variability on short timescales of a few days. On the other hand, a 5120 s low-amplitude and low-significance (2.6-) periodic modulation was detected for NGC 4631 X-6 only in the C2 observation. Hour-scale (4–7 ks) quasi-periodic behavior has been reported in some ULXs (e.g. the M74 ULX; Liu et al. 2005), showing that this timescale is physically plausible for accretion-flow variability. In contrast, orbital periods in ULXs are typically days rather than hours, except for a few systems with Wolf-Rayet donors (e.g, Lin et al., 2023), which have shorter (hour-scale) periods.. Because the period falls below conventional significance thresholds, deeper and higher–signal-to-noise observations are needed to constrain the origin of this periodic feature.
@@
We investigate the relationship between the disc luminosity () and the color temperature of the accretion disc () for the X-6 and X-7. In Galactic BHBs, the X-ray emission in the thermal dominant state is generally produced by a geometrically thin, optically thick disc, which is well described by the standard accretion disc model (Shakura and Sunyaev, 1973). Within this framework, and are expected to follow the relation if the inner disc radius remains constant. Testing this correlation provides a valuable probe of the accretion geometry and physical conditions, while deviations from the expected trend can reveal structural changes in the disc, such as truncation, strong Comptonization, or transitions away from the thin-disc regime, particularly at high luminosities (Kubota and Done, 2004; Done et al., 2007). The Figure 12 shows the relationship between and for ULX X-6 using the spectral parameters of the diskbb model (see Tables 1 and 2). The data reveal a strong positive correlation of the form , with a best-fit slope of (dashed line), consistent within uncertainties with the canonical relation predicted by the standard accretion disk model (Shakura and Sunyaev, 1973). This relation suggests that X-6 is likely in a disk-dominated accretion regime with a relatively stable inner disk radius. The outlier data point (marked in green), denoted C5, significantly deviates from the observed trend, lying at a higher temperature but lower luminosity than expected. This may indicate a temporary transition to a different spectral state, obscuration, or a change in viewing geometry. For instance, Feng and Kaaret (2007) found that NGC 1313 X-2 exhibits a relation consistent with during certain states, indicating a constant inner disk radius. Moreover, transient ULXs such as XMMU J004243.6+412519 in M31 have exhibited L–T tracks broadly consistent with the expected relation. (Middleton et al., 2013). Overall, the trend observed in X-6 reinforces the interpretation of a stellar-mass compact object accreting in a high state, rather than an intermediate-mass BH scenario, which typically exhibits cooler disks at high luminosities (Merloni et al., 2003).
@@
In addition, the normalization parameter of the diskbb model provides an estimate of the apparent inner disk radius through the , where is the normalization, is the observed inner disk radius in km, and is the distance to the source in units of 10 kpc. Using the minimum and maximum normalization values and , the observed inner disk radii are derived 28 and 110 km for minimum and maximum normalization. To account for spectral hardening and boundary condition effects, a correction factor of (Shimura and Takahara, 1995; Kubota et al., 1998) is applied. Therefore, Rin,min and Rin,max are derived as 30 and 130 km. Assuming the inner disk radius corresponds to the innermost stable circular orbit (ISCO) of a Schwarzschild BH, the mass is given by . The estimated mass range is 5–15 , increasing to 5–20 if a disk inclination of 60∘ is assumed. These values are consistent with a stellar-mass BH.
@@
Moreover, we calculated the accretion rate to assess the accretion regime of ULXs following the same approach in Section 4.4 of Ghosh and Rana (2023). We estimate for X-6 an accretion rate of –9 (depending on the assumed compact object mass in the range –). This places the source in a moderately super-Eddington regime, where the spherization radius (Rsph = 270–850 km) lies significantly beyond the inner disk radius Rin = 30–130 km. Such a structure between and is consistent with the expectation that radiation pressure drives strong outflows at larger radii while the inner accretion disk remains relatively compact. The derived beaming factors (–2.3) further suggest that the observed LX of X-6 does not require strong geometric collimation, supporting a scenario in which the source is powered by nearly isotropic emission from a stellar-mass BH in a super-Eddington state (Ghosh and Rana, 2023). In the case of the transient ULX X-7, the vs. relation yields a slope of , which is lower than the () scaling expected for a standard thin disk. However, according to the range (0¡ ¡4) defined by Majumder et al. (2025), X-7 lies within the regime corresponding to a slim-disk behavior (Watarai et al., 2000; Kubota and Makishima, 2004). Similarly, Brightman et al. (2016) found that broadband spectral analyses of ULXs support the presence of optically thick, geometrically slim accretion disks, consistent with the observed parameters of X-7.
4.0.3 Optical counterparts of ULX X-6 and X-7

### v2

@@ 3.2 Short- and long-term X-ray variability
To investigate the short-term X-ray variability of the three newly identified ULXs, we extracted light curves for all available Chandra data in the 0.5–8 keV band and from XMM-Newton data in the 0.3–10 keV band. The analysis was performed using time bins of 100, 500, and 1000 seconds. We used the CIAO and SAS tools dmextract and evselect to background-extract light curves. For the absolute timing analysis, we applied the axbary task to the event file to correct the photon arrival times to the Solar System barycenter, thereby accounting for delays introduced by the orbital motion of the Earth and the Chandra spacecraft.
We applied the Lomb-Scargle periodogram (LS) and fast Fourier transform (FFT) methods to the light curves to search for periodic signals or variability. In the C2 observation of the NGC 4631 ULX X-6 light curve (Figure 4a), we detected a periodic variation at P = 5120 s ( 1.42 hr; 2.6- significance). This signal was identified using both the LS method and with XRONOS from the power density spectrum (Figure 4b). Figure 4c illustrates the folded light curve for this period. To study LX variability, we constructed long-term light curves of X-6 and X-7 from available Chandra data. These luminosities were derived from the best-fit reduced C-stat values obtained in the spectral analysis (Tables 1 and 2). For NGC 1097 ULX-3, we constructed a long-term X-ray light curve using all available archival Swift/XRT PC mode observations, along with data from Chandra and XMM-Newton. We generated the Swift/XRT light curve using a custom Python script (swifttools) designed for Swift/XRT analysis. The source and background regions were circular, with radii of 30′′ and 60′′, respectively. We converted the resulting count rates to LX (0.3-10 keV) using the spectral model parameters from Table 3 and the WebPIMMS tool. For XMM-Newton, we derived luminosities from MOS data based on our spectral modeling, while for Chandra observations, we utilized the srcflux tool. We show the ULX-3 light curve in Figure 6 (left panel).
@@ 3.3 Spectral evolution and state transition
We investigated the spectral evolution of NGC 4631 X-6 and X-7 using Chandra observations, excluding non-detections. Soft (S), medium (M), and hard (H) bands were defined as 0.3–1.2 keV, 1.2–2 keV, and 2–8 keV, respectively, and the M/S and H/M hardness ratios (HR) were calculated. The temporal evolution of the hardness ratios is shown in the left panel of Figure 5, and the hardness–intensity diagram is shown in the right panel. When a band had insufficient counts, we adopted 3 upper limits. To investigate potential spectral state transitions of NGC 1097 ULX-3, we constructed a hardness–intensity diagram (right panel of Figure 6), using count rates in the 0.3–1.5 keV (S) and 1.5–10 keV (H) bands. To account for differences in instrumental responses, we normalized the count rates from Chandra and XMM-Newton observations to those from Swift/XRT using WebPIMMS.
@@ 3.4 Optical and infrared photometry
For all drizzled HST/Wide Field Camera 3/Ultraviolet and Visible channel (WFC3/UVIS) and HST/Advanced Camera for Surveys/Wide Field Channel (ACS/WFC) images, we identified point-like sources with daofind task, and their fluxes were measured with aperture photometry employing the DAOPHOT package (Stetson, 1987) within IRAF222https://iraf-community.github.io/ (Image Reduction and Analysis Facility). For the JWST/NIRCam observations, background estimation, source detection, and photometry were performed following the methodology of Allak (2023). Instrumental magnitudes were converted into Vega magnitudes using zero points (ZP) provided for WFC3/UVIS and ACS/WFC taken from Calamida et al. (2022), and the ZP calculator333https://acszeropoints.stsci.edu/. Sources were required to exceed a 3 significance above the local background. We then measured photometry using circular apertures of 3-pixel radius, with the background contribution subtracted using an annulus located nine pixels from the centroid.
Using the astrometric corrections reported by Guo et al. (2023), we identified three optical counterparts (X6a-Xc) for X-6 and two (X7a and X7b) for X-7 in the F606W and F814W bands. For NGC 1097 ULX-3, we searched the available catalogs (2MASS and Gaia DR3) for optical and NIR counterparts. Aside from the galaxy nucleus, we did not identify any reference sources with XRBs in these catalogs; therefore, the nucleus was used as the reference source for astrometric corrections. We corrected the relative astrometry between Chandra (CH7) and HST (F555W), Chandra and JWST (F115W), and HST and JWST, obtaining confidence error radii of , , and , respectively. Based on these results, we identified a unique optical and NIR counterpart within the astrometric error radius by applying a 3 detection threshold over the background. Figure 7 shows the positions of the ULXs in Pan-STARRS RGB composite images (i, r, g bands), illustrating their distribution across the galaxy. In the NIRCam images (Figure 8), several faint NIR sources are present around the counterpart, but were not detected above the 1.5 level. Therefore, we adopted the optical counterpart as the corresponding NIR counterpart of ULX-3. Figure 8 shows the positions of all three ULX counterparts. Tables 4 and 5 list the photometric results for the optical and NIR counterparts of NGC 4631 ULX X-6 and X-7, and NGC 109 ULX-3.
@@
Color–magnitude diagrams (CMDs) provide insights into the ages of the stellar populations surrounding ULXs, which may help constrain the nature of their donor stars. Isochrone models from the PAdova TRieste Stellar Evolution Code (PARSEC) (Bressan et al., 2012) stellar evolution code are available for various telescope filters, including those of HST/ACS. We derived the upper limit extinction, AV, using the method of Guo et al. (2023) and used this value to correct the isochrones. Using the average NH values from Tables 1 and 2, we derived E(F606W-F814W) values of 2.90 mag and 1.72 mag for X-6 and X-7, respectively. Figure 9 shows the CMDs for X-6 and X-7, plotted as F606W versus F606W-F814W, with isochrones and field sources located within 2′′ radius circles centered on the ULX positions, alongside the optical counterparts. For ULX-3, we only constructed an NIR CMD due to the limited optical data. The HST/F814W photometry shows variability likely originating from the accretion disk; nonsimultaneous filters prevent reliable color measurements. The source is extremely faint in the UVIS bands and F438W detection is consistent with an upper limit, making an optical CMD infeasible. In contrast, the JWST/NIRCam data show no variability; therefore, we were able to construct a single-epoch NIR CMD. The NIR CMD was constructed using F115W and F277W photometry, adopting a distance of 16.8 Mpc and an upper limit extinction of 0.06 mag (Figure 10).
For ULX-3, we converted the observed magnitudes into fluxes using the corresponding zeropoints and adopted the pivot wavelength of each JWST/NIRCam filter. The resulting SED is well fit by a blackbody model with a best-fit temperature of T = 3300 140 K and a reduced chi–square of = 0.95 for four degrees of freedom (Figure 11).
4 Results and discussion
4.1 NGC 4631 X-6 and X-7
@@ 4.1.1 Spectral modeling and variability
In NGC 4631, we identified two new transient ULX candidates, X-6 and X-7, from observations conducted between 2021 and 2023. Across 13 Chandra observations (C1–C13) and three XMM-Newton observations (XM1–XM3), we did not detect X-6 in C1, XM1, or XM3. We detected X-7 source only in XM2 and in five Chandra observations (C2–C7). We also examined Swift/XRT PC observations using the Swift-XRT data products tool444https://www.swift.ac.uk/user_objects/ to investigate the visibility of X-6 and X-7.
@@
For X-6, both the power-law and diskbb models yield values that vary by nearly an order of magnitude in Chandra observations when intrinsic absorption is left free. The variability may reflect genuine changes in the local environment, but could also result from statistical uncertainties or spectral modeling limitations. We therefore fixed to its average value and repeated the spectral fits (Table 1). After applying this procedure, we obtained notable improvements in the reduced values for both the diskbb and power-law models. Applying simple phenomenological models to the X-6 spectra, we found that the two models could not be distinguished in the Chandra datasets, likely because of limited data quality. The power-law photon index () ranges between 1.6 and 2.8. For Galactic BH binaries (BHBs), 1.4–2.4 typically corresponds to the hard state, characterized by Comptonization of disk photons in a hot, optically thin corona (e.g., Remillard and McClintock, 2006; Done et al., 2007). For X-6, remains ¡2.4 in nearly all Chandra observations (except C8), indicating that the source resides predominantly in the hard state. Based on the power-law model, X-6 exhibits LX = erg s-1 in the 0.5–8 keV band.
@@
Furthermore, X-6 is well described by the diskbb model, with the inner disk temperature () varying between 0.8 and 2 keV. Using the diskbb spectral model, X-6 exhibits LX = erg s-1 in the 0.5–8 keV band. For X-7, among all the single and multi-component models tested in XSPEC, only the diskbb model provides acceptable fits, yielding inner disk temperatures of – keV for roughly half of the dataset. For the C2–C7 observations, X-7 exhibits LX = erg s-1 in the 0.5–8 keV band, based on the diskbb spectral model. The temperatures indicate that X-ray emission is dominated by accretion disk processes, consistent with Galactic XRBs in the thermal state at high accretion rates. Such disk-dominated spectral behavior has been reported in several ULX studies, with thermal components consistent with super-Eddington accretion or slim disk models (e.g., Gladstone et al., 2009; Luangtip et al., 2016). To search for spectral state transitions, we constructed hardness ratios as a function of time and hardness–intensity diagrams (Figure 5). Source X-6 predominantly occupies the hard state, with episodes indicating transitions from hard to intermediate and back to hard, whereas source X-7 resides mostly in the intermediate state.
@@
The X-6 and X-7 ULXs do not show significant variability on short timescales of a few days. We detected a 5120 s, low-amplitude, low-significance (2.6-) periodic modulation for NGC 4631 X-6 only in the C2 observation. Hour-scale (4–7 ks) quasi-periodic behavior has been reported in some ULXs (e.g., the M74 ULX; Liu et al. 2005), showing that this timescale is physically plausible for accretion-flow variability. In contrast, orbital periods in ULXs are typically days, except for a few systems with Wolf-Rayet donors (e.g., Lin et al., 2023), which show hour-scale periods. Because the period falls below conventional significance thresholds, deeper and higher signal-to-noise observations are needed to constrain the origin of this periodic feature.
@@
We investigated the relationship between disk luminosity () and the color temperature of the accretion disk () for X-6 and X-7. In Galactic BHBs, the X-ray emission in the thermal dominant state is generally produced by a geometrically thin, optically thick disk, which is well described by the standard accretion disk model (Shakura and Sunyaev, 1973). Within this framework, and are expected to follow the relation if the inner disk radius remains constant. Testing this correlation probes accretion geometry and physical conditions, while deviations from the expected trend can reveal structural changes in the disk, such as truncation, strong Comptonization, or transitions away from the thin-disk regime, particularly at high luminosities (Kubota and Done, 2004; Done et al., 2007). Figure 12 shows the relationship between and for X-6 using diskbb spectral parameters (Tables 1 and 2). The data reveal a strong positive correlation of the form , with a best-fit slope (dashed line), consistent within the uncertainties with the canonical relation predicted by the standard accretion disk model (Shakura and Sunyaev, 1973). This relation suggests that X-6 is likely in a disk-dominated accretion regime with a relatively stable inner disk radius. The outlier, denoted C5, significantly deviates from the observed trend, lying at a higher temperature but lower luminosity than expected. This may indicate a temporary transition to a different spectral state, obscuration, or a change in viewing geometry. For example, Feng and Kaaret (2007) found that NGC 1313 X-2 exhibited a relation consistent with during certain states, indicating a constant inner disk radius. Transient ULXs, such as XMMU J004243.6+412519 in M31, have also shown L–T tracks broadly consistent with the expected relation. (Middleton et al., 2013). Overall, the trend observed in X-6 supports a stellar-mass compact object accreting in a high state, rather than an intermediate-mass BH scenario, which typically exhibits cooler disks at high luminosities (Merloni et al., 2003).
@@
In addition, the diskbb normalization parameter provides an estimate of the apparent inner disk radius: , where is the normalization, is the observed inner disk radius in km, and is the distance to the source in units of 10 kpc. Using and , the observed inner disk radii are 28 and 110 km. To account for spectral hardening and boundary condition effects, we applied a correction factor of (Shimura and Takahara, 1995; Kubota et al., 1998). Therefore, the corrected values of Rin,min and Rin,max are 30 and 130 km, respectively. Assuming that the inner disk radius corresponds to the innermost stable circular orbit (ISCO) of a Schwarzschild BH, , the mass is . The estimated mass range is 5–15 , increasing to 5–20 for an assumed disk inclination of 60∘. These values are consistent with a stellar-mass BH.
@@
We calculated the accretion rate to assess the accretion regime of ULXs following the method described in Section 4.4 of Ghosh and Rana (2023). For X-6, we estimate an accretion rate of –9, depending on the assumed compact object mass in the range –. This estimate places the source in a moderately super-Eddington regime, where the spherization radius (Rsph = 270–850 km) lies significantly beyond the inner disk radius, Rin = 30–130 km. This structure between and is consistent with the expectation that radiation pressure drives strong outflows at larger radii, while the inner accretion disk remains relatively compact. The derived beaming factors (–2.3) further suggest that the observed LX of X-6 does not require strong geometric collimation, supporting a scenario in which the source is powered by nearly isotropic emission from a stellar-mass BH in a super-Eddington state (Ghosh and Rana, 2023). For the transient ULX X-7, the versus relation yields a slope of , which is lower than the () scaling expected for a standard thin disk. However, according to the range (0¡ ¡4) defined by Majumder et al. (2025), X-7 lies within the regime corresponding to slim-disk behavior (Watarai et al., 2000; Kubota and Makishima, 2004). Similarly, Brightman et al. (2016) report that broadband spectral analyses of ULXs support the presence of optically thick, geometrically slim accretion disks, consistent with the observed parameters of X-7.
4.1.3 Optical counterparts of ULXs X-6 and X-7